# Bias Amplification: Large Language Models as Increasingly Biased Media

### v2

@@ Abstract
Model collapse—a phenomenon where models degrade in performance due to indiscriminate use of synthetic data—is well studied. However, its role in bias amplification—the progressive reinforcement of pre-existing social biases in Large Language Models (LLMs)—remains underexplored. In this paper, we formally define the conditions for bias amplification and demonstrate through statistical simulations that bias can intensify even in the absence of sampling errors, the primary driver of model collapse. Empirically, we investigate political bias amplification in GPT-2 using a custom-built benchmark for sentence continuation tasks. Our findings reveal a progressively increasing right-leaning bias. Furthermore, we evaluate three mitigation strategies—Overfitting, Preservation, and Accumulation—and show that bias amplification persists even when model collapse is mitigated. Finally, a mechanistic interpretation identifies distinct sets of neurons responsible for model collapse and bias amplification, suggesting they arise from different underlying mechanisms.
@@ 1 Introduction
Large Language models (LLMs) are trained on vast amounts of text scraped from the internet, which plays a crucial role in improving their capabilities, whether through emergent abilities (Wei et al., 2022) or scaling laws (Kaplan et al., 2020). However, as they become more widely integrated into human society—for example, in content creation and summarization in media, academia, and business (Maslej et al., 2024)—concerns are mounting that a significant portion of online text in the future may be generated, either entirely or partially, by LLMs (Peña-Fernández et al., 2023; Porlezza and Ferri, 2022; Nishal and Diakopoulos, 2024). This highlights a significant and underexplored risk: bias amplification, where pre-existing biases become progressively reinforced and intensified as models are repeatedly trained on synthetic data (Mehrabi et al., 2022; Taori and Hashimoto, 2022). This concern initially arises from the tendency of LLMs to learn from biased datasets. For example, Parrish et al. (2022); Wang et al. (2024); Bender et al. (2021) show that LLMs absorb inherent biases embedded in human-generated text. Haller et al. (2023); Rettenberger et al. (2024a) demonstrated that LLMs can be aligned with specific political ideologies by fine-tuning them on biased datasets. Moreover, Wyllie et al. (2024) shows that classifiers trained on synthetic data increasingly favor a particular class label over successive generations, while the shrinking diversity observed by Alemohammad et al. (2023); Hamilton (2024) suggests a risk that certain demographic groups may become underrepresented in the outputs of LLMs.
The amplification of biases has profound societal implications. It can lead to the perpetuation of stereotypes, reinforcement of social inequalities, and the marginalization of underrepresented groups. In the context of political bias, this can influence public opinion, skew democratic processes, and exacerbate polarization. Therefore, understanding and mitigating bias amplification is both critical and urgent. Nevertheless, despite the literature on discriminative models, there is a notable lack of comprehensive frameworks and empirical studies specifically addressing bias amplification in Language Model.
Figure 1: Experimental Procedure.
In this paper, we aim to address this research gap by introducing a framework that explains the underlying causes of bias amplification in LLMs. We validate the framework through both statistical simulations and direct experiments, demonstrating the emergence of bias amplification. Furthermore, our findings show the potential distinction between bias amplification and model collapse. In summary, our key contributions are as follows:
1. Formal Conditions: We formally delineate the conditions for bias amplification (Section 3.1), providing the theoretical basis for understanding its causes and relationship to model collapse. We then validate this framework through statistical simulations leveraging a weighted maximum likelihood estimation cycle (Section 3.2).
2. Political Bias Benchmark: We trained a highly accurate classifier capable of detecting political leaning in long-text content, specifically within the context of the US political spectrum. With this classifier as a metric model, we offer a benchmark for evaluating political bias in LLMs through sentence continuation tasks.
3. Empirical Validation: We evaluate the amplification of political bias in GPT-2, where the model exhibits a right-leaning bias in sentence continuation tasks and shifts further to the right over generations (Section 5.1). Additionally, we experimented with three potential mitigation strategies—Overfitting, Preservation, and Accumulation—comparing their effectiveness in both Model collapse and Bias Amplification (Section 5.3). This experiment setup could be easily extended to larger models or different types of bias.
4. Mechanistic Interpretation: We conduct a mechanistic analysis identifying two distinct sets of neurons responsible for bias amplification and model collapse during iterative synthetic fine-tuning with GPT-2. Our findings reveal only partial overlap between these sets, suggesting that bias amplification and model collapse arise from distinct mechanisms (Section 5.4).
2 Background and Related Work
Bias Amplification has been studied in various domains Zhao et al. (2017) found that Conditional Random Fields can exacerbate social biases present in the training data. It proposed an in-process mitigation approach, employing Lagrangian Relaxation to enforce constraints that ensure the model’s bias performance remains closely aligned with the biases in the training data. Following this, Mehrabi et al. (2022) proposed the concept of bias amplification in feedback loops, where biased models not only amplify the bias present in their training data but also interact with the world in ways that generate more biased data for future models. Xu et al. (2023); Zhou et al. (2024) examined bias amplification in recommendation models, showing how these models reinforce their understanding of mainstream user preferences from training data, leading to an overrepresentation of such preferences in historical data and neglecting rarely exposed items—similar to the concept of sampling error discussed in (Shumailov et al., 2024). Wyllie et al. (2024); Taori and Hashimoto (2022) demonstrated that classifiers trained on synthetic data increasingly favor specific class labels over successive generations. Likewise, Ferbach et al. (2024); Chen et al. (2024) observed bias amplification in generative models such as Stable Diffusion, characterized by the overrepresentation of features from the training dataset. Also, we provide a comprehensive literature review of model collapse in Appendix J.
(a) Weighted Maximum Likelihood Estimation.
(b) Maximum Likelihood Estimation.
Figure 2: Comparison of WMLE and MLE over 10 generations.
@@ 2 Background and Related Work
Political Biases. In parallel, growing attention has been paid to political biases in LLMs, now a prevalent form of “media” that people rely on for global news (Maslej et al., 2024). Rettenberger et al. (2024b); Shumailov et al. (2024); Feng et al. (2024) explored the bias through voting simulations within the spectrum of German political parties, consistently finding a left-leaning bias in models like GPT-3 and Llama3-70B. Similarly, for the U.S. political landscape, Rotaru et al. (2024); Motoki et al. (2024) identified a noticeable left-leaning bias in ChatGPT and Gemini when tasked with rating news content, evaluating sources, or responding to political questionnaires.
3 Formal Conditions
In this section, we formalize the conditions for bias amplification, offering an intuitive look at its principal drivers. We then illustrate these ideas using Weighted Maximum Likelihood Estimation (WMLE).
3.1 The Conditions for Bias Amplification
The primary factor is referred to as bias projection. It arises when the bias projection coefficient is negative. To illustrate this, consider a fine-tuning process in which the pre-trained model parameters $\theta_{t}$ can be expressed as the sum of unbiased and biased components:
$\theta_{t}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}.$
During gradient-based optimization, the update rule is:
If $c_{t}<0$, the gradient update will reinforce the biased component, leading to bias amplification, i.e. $\Delta|\theta_{\text{biased}}|>0$. This occurs because the gradient descent step moves the parameters further in the direction of the existing bias. Intuitively, bias in the output arises from specific neurons, whose weights increase when their activations align with the fine-tuning dataset. If biased neurons capture more relevant patterns than unbiased ones, optimization will further reinforce their weights.
The second is sampling error, akin to statistical approximation error Shumailov et al. (2024). If the model has a pre-existing bias, it inherently assigns higher probabilities to tokens that produce biased outputs. Consequently, during synthetic data generation, unbiased tokens—and thus unbiased samples—are more likely to be lost at each resampling step with a finite sample, though this error vanishes as the sample size approaches infinity. This overrepresents biased patterns in the synthetic data, surpassing the model’s original bias and true next-token probabilities. Sampling error thus complements bias projection by further activating biased neurons in response to the skewed dataset.
By definition, bias projection is a sufficient condition for bias amplification, while sampling error serves as a complementary factor. However, sampling error is a sufficient condition for model collapse to occur with nonzero probability Shumailov et al. (2024). This distinction motivates our investigation into whether bias amplification can occur without model collapse, as this would necessitate additional mitigation strategies beyond those for model collapse.
3.2 Statistical Simulation
To simulate a controlled setting without sampling error, we consider a statistical estimation cycle using WMLE [^1] The mathematical formulation is detailed in Appendix A. with a large sample size of each resampling step. Specifically, we generate a pre-training dataset $\mathcal{D}_{\text{pre}}$ with 100,000 samples from a Beta$(3,2)$ distribution, representing a biased pre-training dataset. Using maximum likelihood estimation (MLE), we estimate its probability density function, yielding the pre-trained model $f_{\text{pre}}$.
Figure 2(a) shows the estimated distributions gradually shift toward the mean of the biased pre-training dataset at $x=0.6$, becoming progressively more peaked over generations. This occurs despite further training on samples drawn from Beta(2,2)22(2,2)( 2 , 2 ) and synthetic data generated from successive models. The distortion arises because the fine-tuning process disproportionately emphasizes regions where the pre-trained distribution assigns higher probability, leading to biased learning. For comparison, Figure 2(b) presents the results using standard MLE without weighting. In this case, the estimated distributions remain stable across generations, accurately representing the Beta(2,2)22(2,2)( 2 , 2 ) distribution.
4 Experimental Design
@@
This section provides the details of the experiments on LLMs, focusing on the sequential and synthetic fine-tuning of GPT-2 under different setups [^2] Similar experiments with larger models like GPT-4 or Claude can be easily adapted to the current setup but would incur significant environmental costs.. The step-by-step experimental procedure is outlined in Figure 1. Our study focuses on the political bias of LLMs within the US political spectrum, particularly in sentence continuation tasks. This is important as LLMs are increasingly influencing global news consumption Maslej et al. (2024); Peña-Fernández et al. (2023); Porlezza and Ferri (2022), and traditional news outlets, such as the Associated Press, are beginning to integrate LLMs for automated content generation from structured data (The Associated Press, 2024).
@@ 4.2 Successive Fine-tuning
Building on the approach in Shumailov et al. (2024); Dohmatob et al. (2024b), each training cycle begins with fine-tuning GPT-2 on a dataset of 1,518 news articles. The fine-tuned model, referred to as Generation 0, generates a synthetic dataset $D_{0}$ of the same size, which is then used for fine-tuning to produce Generation 1. This iterative process continues until Generation 10 is reached. Note that synthetic fine-tuning starts at Generation 1.
@@ 4.3 Synthetic Data Generation
Synthetic datasets, $\{D_{i}\}_{i=0}^{10}$, are generated as follows: Each original news article is tokenized into 64-token blocks, and for each block, the model predicts the next 64 tokens using deterministic generation to enhance the replicability of the study. The generated tokens are then decoded back into text, producing a synthetic dataset of the same size as the original.
@@ 4.4 Political Bias Metric
We develop a classification model to assess the political leaning of each LLM based on its generated synthetic news articles. The model is trained on the Webis-Bias-Flipper-18 dataset, excluding the 1,518 articles used for GPT-2 fine-tuning. To mitigate class imbalance, center-leaning articles are resampled to ensure equal representation across categories. The dataset is then divided into training (70%), validation (15%), and test (15%) subsets, stratified by bias label. Additionally, we conduct a human review to remove any identifiable information about media sources and authors.
After performing a grid search across multiple models, we find that roberta-base achieves the best performance, with an evaluation loss of 0.4035 and a macro F1 score of 0.9196 on the test set (see Table 1). Thus, we select roberta-base as the benchmark for political bias detection in subsequent experiments. Details on model training are provided in Appendix C.
Table 1: Evaluation Results for Political Bias Classifier
@@ 4.5 Generation Quality Metric
We introduce a metric to evaluate generation quality, specifically addressing the issue of repetitive content in later model iterations, which can distort traditional perplexity metrics. This metric is based on the Gibberish Detector (Jindal, 2021), which identifies incoherent or nonsensical text. The detector categorizes text into four levels: (1) Noise—individual words hold no meaning, (2) Word Salad—incoherent phrases, (3) Mild Gibberish—grammatical or syntactical distortions, and (4) Clean—coherent, meaningful sentences.
To quantify generation quality, each sentence receives a Gibberish score: 3 for Clean, 2 for Mild Gibberish, 1 for Word Salad, and 0 for Noise. The text quality index is computed as the average score across all sentences in an article. This metric prioritizes coherence and meaning, offering a better assessment of generation quality than perplexity.

### v3

@@ Abstract
Model collapse—a phenomenon characterized by performance degradation due to iterative training on synthetic data—has been widely studied. However, its implications for bias amplification, the progressive intensification of pre-existing societal biases in Large Language Models (LLMs), remain significantly underexplored, despite the growing influence of LLMs in shaping online discourse. In this paper, we introduce a open, generational, and long-context benchmark specifically designed to measure political bias amplification in LLMs, leveraging sentence continuation tasks derived from a comprehensive dataset of U.S. political news. Our empirical study using GPT-2 reveals consistent and substantial political bias intensification (e.g., right-leaning amplification) over iterative synthetic training cycles. We evaluate three mitigation strategies—Overfitting, Preservation, and Accumulation—and demonstrate that bias amplification persists independently of model collapse, even when the latter is effectively controlled. Furthermore, we propose a mechanistic analysis approach that identifies neurons correlated with specific phenomena during inference through regression and statistical tests. This analysis uncovers largely distinct neuron populations driving bias amplification and model collapse, underscoring fundamentally different underlying mechanisms. Finally, we supplement our empirical findings with theoretical intuition that explains the separate origins of these phenomena, guiding targeted strategies for bias mitigation.
@@ 1 Introduction
Large Language Models (LLMs) have become essential tools for content creation and summarization in various sectors, including media, academia, and business (Maslej et al., 2024). However, a significant but underexplored risk arises as LLMs increasingly rely on their own or other synthetic outputs for training, potentially amplifying pre-existing societal biases (Peña-Fernández et al., 2023; Porlezza and Ferri, 2022; Nishal and Diakopoulos, 2024). This phenomenon, known as bias amplification, refers to the progressive reinforcement and intensification of existing biases through iterative synthetic training (Mehrabi et al., 2022; Taori and Hashimoto, 2022). This issue stems from the inherent tendency of LLMs to learn from biased datasets; prior studies indicate that LLMs readily absorb biases from human-generated text (Parrish et al., 2022; Wang et al., 2024; Bender et al., 2021). Further, fine-tuning LLMs on biased datasets can align them with specific political ideologies (Haller et al., 2023; Rettenberger et al., 2024a). Classifiers trained on synthetic data also increasingly favor specific class labels over time (Wyllie et al., 2024), and diversity tends to decrease, potentially marginalizing certain demographic groups (Alemohammad et al., 2023; Hamilton, 2024). The implications of bias amplification are substantial, including perpetuation of stereotypes, reinforcement of social inequalities, and potential impacts on democratic processes through the skewing of public opinion and increased polarization. Despite its significance, comprehensive frameworks and empirical research specifically addressing bias amplification in language models remain sparse, although related work exists on discriminative models.
Figure 1: Overview of the iterative experimental procedure for synthetic fine-tuning and analysis.
In this study, we empirically investigate political bias amplification in GPT-2. We define political bias as the disproportionate generation of content aligned with specific political ideologies. Our experiments reveal that GPT-2 progressively exhibits stronger right-leaning and center-leaning biases in two distinct scenarios: (1) starting from fine-tuning on an unbiased dataset, and (2) initially fine-tuned exclusively on center-leaning articles. We also evaluate three mitigation strategies—Overfitting, Preservation, and Accumulation—to address bias amplification and model collapse. Preservation mitigates both phenomena effectively in the first scenario but fails to prevent bias amplification in the second. Additionally, we propose a novel mechanistic analysis using regression and statistical testing to examine neuron-level changes correlated with bias amplification and model collapse, identifying largely distinct neuron groups for each phenomenon. This suggests different underlying mechanisms for bias amplification and model collapse.
In summary, our contributions are: (i) a highly accurate classifier for detecting political bias in long-text content, providing a benchmark for evaluating political bias in LLMs via sentence continuation tasks; (ii) an empirical assessment of political bias amplification in GPT-2 across two fine-tuning setups; (iii) evaluation of three mitigation strategies; (iv) a novel mechanistic analysis method identifying neurons correlated with specific phenomena during inference; and (v) a theoretical intuition that explains the difference between bias amplification and model collapse based on their underlying causes. The experimental framework presented can be extended to other models and different types of biases.
2 Related Work
Bias Amplification has been studied in various domains. For instance, Zhao et al. (2017) found Conditional Random Fields can worsen social biases from training data, proposing an in-process Lagrangian Relaxation method to align model and data biases. Mehrabi et al. (2022) later described bias amplification in feedback loops, where models amplify existing bias and generate more biased data through real-world interaction. Xu et al. (2023); Zhou et al. (2024) showed recommendation models amplify mainstream preferences, overrepresenting them and neglecting rarer items, akin to sampling error (Shumailov et al., 2024).
Classifiers trained on synthetic data increasingly favor certain labels over generations (Wyllie et al., 2024; Taori and Hashimoto, 2022). Similarly, generative models like Stable Diffusion show bias amplification through feature overrepresentation from training data (Ferbach et al., 2024; Chen et al., 2024). More recently, Li et al. (2025) investigated gender and cultural bias amplification in LLMs (classification and generation tasks, 1-5 synthetic rounds), proposing pre-processing (labeling bias, removing identity words) and in-processing (penalizing deviation from real data) mitigations; these showed varied effectiveness in one-round fine-tuning.
Model Collapse. Model collapse is a deterioration where models recursively trained on their own output distort reality and lose generalizability (e.g., prioritizing common events, neglecting rare ones, or shifting distributions) (Shumailov et al., 2024; Alemohammad et al., 2023; Guo et al., 2024; Wyllie et al., 2024; Dohmatob et al., 2024a). Shumailov et al. (2024) showed this with OPT-125M, where perplexity distributions skewed towards lower values with longer tails. Increased repetition in synthetically fine-tuned GPT-2 was noted by Taori and Hashimoto (2022). Performance deterioration in models like OPT-350M, Llama2, and GPT-2 (e.g., reduced linguistic diversity, token probability divergence) after several generations was shown by Guo et al. (2024); Dohmatob et al. (2024b); Seddik et al. (2024). In generative image models, Alemohammad et al. (2023) found quality and diversity deteriorate with synthetic training; however, user cherry-picking of high-quality outputs (a form of sampling error) helped maintain quality. Hamilton (2024) noted GPT-3.5-turbo exhibited less perspective diversity in narrative writing than earlier models (davinci-instruct-beta, text-davinci-003).
@@ 2 Related Work
Political Biases. In parallel, growing attention has been paid to political biases in LLMs, now a prevalent form of "media" that people rely on for global news (Maslej et al., 2024). Rettenberger et al. (2024b); Shumailov et al. (2024); Feng et al. (2024) explored the bias through voting simulations within the spectrum of German political parties, consistently finding a left-leaning bias in models like GPT-3 and Llama3-70B. Similarly, for the U.S. political landscape, Rotaru et al. (2024); Motoki et al. (2024) identified a noticeable left-leaning bias in ChatGPT and Gemini when tasked with rating news content, evaluating sources, or responding to political questionnaires. (Bang et al., 2024) study political bias in LLMs through the task of generating news headlines on politically sensitive topics and find that the political perspectives expressed by LLMs vary depending on the subject matter.
3 Methodology
@@
This section provides the details of the experiments on LLMs, focusing on the sequential and synthetic fine-tuning of GPT-2. The step-by-step experimental procedure is outlined in Figure 1. Our study focuses on the political bias of LLMs within the US political spectrum, particularly in sentence continuation tasks. This is important as LLMs are increasingly influencing global news consumption Maslej et al. (2024); Peña-Fernández et al. (2023); Porlezza and Ferri (2022), and traditional news outlets, such as the Associated Press, are beginning to integrate LLMs for automated content generation from structured data (The Associated Press, 2024).
@@ 3.2 Successive Fine-tuning
[…]
The fine-tuning procedure remained consistent across all experiments unless stated otherwise. The input length was capped at 512 tokens, with the EOS token used for padding. The model was trained for 5 epochs, using a batch size of 8, a learning rate of $5\times 10^{-5}$, and a weight decay of 0.01. Fine-tuning was conducted using standard functionalities available in transformer libraries. After each cycle, the model was saved and used to generate synthetic data for the subsequent iteration. [^1] Fine-tuned models will be made public upon acceptance.
@@ 3.3 Synthetic Data Generation
Synthetic datasets, $\{D_{i}\}_{i=0}^{10}$, are generated as follows: For each original news article, its tokenized title serves as an initial prompt, and its tokenized body is segmented into sequential 64-token blocks, which serve as subsequent prompts. For each such prompt, the model predicts the next 64 tokens. These predictions are made using deterministic generation to enhance the reproducibility. All the newly generated 64-token sequences (one from the title prompt and one from each body block prompt) are concatenated and then decoded back into text. This process creates one synthetic article from each original article, resulting in a synthetic dataset of the same number of articles as the original.
@@ 3.4 Political Bias Metric
Metric. We develop a classification model to assess the political leaning of each LLM based on its generated synthetic news articles. The model is trained on the Webis-Bias-Flipper-18 dataset, excluding the 1,518 articles used for GPT-2 fine-tuning. To mitigate class imbalance, center-leaning articles are resampled to ensure equal representation across categories. The dataset is then divided into training (70%), validation (15%), and test (15%) subsets, stratified by bias label. Additionally, we conduct a human review to remove any identifiable information about media sources and authors.
Specifically, the training dataset comprises 2,781 distinct events that occurred in the U.S. between 2012 and 2018. For each event, it includes news articles collected from a wide range of media outlets. In total, it contains articles from 97 different outlets, such as The Washington Examiner, The Washington Post, HuffPost, Reuters, and others. This makes the dataset a strong representation of the diversity of U.S. media sources and event domains, and therefore strengthens the generalizability of our classifier in the U.S. context.
After performing a grid search across multiple models, we find that roberta-base achieves the best performance, with an evaluation loss of 0.4035 and a macro F1 score of 0.9196 on the test set (see Table 1). Thus, we select roberta-base as the benchmark for political bias detection in subsequent experiments. It is important to note that this classifier is trained on data from 2012-2018, and its direct application to significantly later content might eventually require recalibration due to the evolving nature of the political landscape and discourse. Details on model training are provided in Appendix A.
Bias Construct. We define political bias in the news article generation task as the disproportionate production of articles with specific political leanings, as identified by our classifier. Unlike Bang et al. (2024), which defines political bias in a topic- or issue-specific manner, we take a broader perspective by measuring the overall political leaning of the model across a diverse set of topics. This approach uses articles from U.S. media outlets published between 2012 and 2018. We use this definition because we view LLMs as analogous to media outlets: while an outlet may publish content spanning the political spectrum, both the public and bipartisan organizations like AllSides assign it an overall political rating based on the average ideological slant of its content and its perceived alignment.
Table 1: Macro F1 Scores for Political Bias Classifier Models; roberta-base selected.
@@ 3.5 Generation Quality Metric
Metric. We introduce a metric, text quality index, to evaluate generation quality, specifically addressing the issue of repetitive content in later model iterations, which can distort traditional perplexity metrics (see Section 4.2). This metric is based on the Gibberish Detector (Jindal, 2021), which identifies incoherent or nonsensical text. The detector categorizes text into four levels: (1) Noise—individual words hold no meaning, (2) Word Salad—incoherent phrases, (3) Mild Gibberish—grammatical or syntactical distortions, and (4) Clean—coherent, meaningful sentences. To quantify generation quality, each sentence receives a Gibberish score: 3 for Clean, 2 for Mild Gibberish, 1 for Word Salad, and 0 for Noise. The text quality index is computed as the average score across all sentences in an article. This metric prioritizes coherence and meaning, offering a more meaningful assessment of generation quality than perplexity.
Motivation. We adopt the gibberish detector as our primary tool for evaluating generation quality because it directly captures a fundamental aspect of model deterioration: the loss of coherence and semantic clarity, which becomes particularly pronounced in later generations of synthetic training as observed in our experiments (see Section 4). While other forms of degradation—such as reduced factual consistency or increased hallucination—may also occur, they are often secondary to the loss of basic intelligibility. As such, we leave the investigation of factual consistency deterioration to future work, particularly in contexts more narrowly focused on diagnosing model collapse.
3.6 Mechanistic Analysis
To gain a clearer understanding of the causes of bias amplification and how it empirically relates to model collapse, we conduct a mechanistic analysis of how neurons behave and vary across different generations [^2] We have 11 generations for each training round, with a total of 6 rounds, resulting in 66 versions of fine-tuned GPT-2. of fine-tuned GPT-2 models, each exhibiting different levels of generation quality and bias performance.
The first step is to extract the changing weight (or activation value) pattern of each neuron across versions and compare it to the corresponding changes in bias performance and generation quality. For each of the 9,216 neurons, which correspond to the 768 output neurons from the feed-forward network (FFN) sublayer in each of the 12 transformer blocks of the GPT-2 model, we compute the correlation between its weight (or activation value) and the model’s bias performance (or generation quality) across all 66 versions.
Figure 2: Distribution of political bias labels (’Left’, ’Center’, ’Right’) for initial GPT-2 synthetic outputs, classified by our Political Bias Metric.
To statistically test the significance of these correlations, we estimate the following linear model:
$\Delta y_{i}=\alpha_{j}+\beta_{j}\Delta x_{i,j}+\epsilon_{i,j}$ (1)
where $\Delta y_{i}$ denotes the change in the proportion of articles leaning in a specific political direction (e.g., the proportion of right-leaning articles if the model is biased in that direction), or the change in the text quality index, between model $i-1$ and model $i$. The term $\Delta x_{i,j}$ represents the change in the weight (or activation value) of neuron $j$ over the same transition. The coefficient $\beta_{j}$ captures the extent to which changes in the weight (or activation value) of neuron $j$ are associated with shifts in political bias (or generation quality), while $\alpha_{j}$ is a constant and $\epsilon_{i,j}$ is the residual error. By applying first-order differencing to both $x_{i,j}$ and $y_{i}$, we reduce potential serial correlation, ensuring that our regression estimates better reflect the dynamic influence of individual neuron weight updates.
We assess the statistical significance of each $\beta_{j}$ using Newey-West adjusted p-values. [^3] Details for the statistical tests are provided in Appendix G. Using the $p$-values and a $95\%$ significance threshold, we identify the sets of neurons significantly correlated with bias amplification (i.e., changes in the proportion of politically leaning articles) and with model collapse (i.e., changes in the generation quality index). By comparing these sets and analyzing their degree of overlap, we gain evidence about whether
(a) Right-leaning bias %
(b) Text quality index
Figure 3: Evolution of (a) right-leaning bias and (b) text quality index across generations (initial G0: unbiased dataset). Compares baseline (’Synthetic’) with three mitigation strategies. Text quality includes 95% CIs.
the two phenomena arise from distinct underlying mechanisms.